OpenStudy (anonymous):

- For a substance C, the time rate of conversion is proportional to the square of the amount x of unconverted substance. Let k be the numerical value of the constant of proportionality and let the amount of unconverted substance be x0 at time t=0. Determine x for all t (greater than or equal to) 0.

OpenStudy (dumbcow):

wait i am wrong...this is a differential equation problem huh? \[\frac{dx}{dt} = kx^{2}\]

OpenStudy (dumbcow):

separate variables \[\frac{dx}{x^{2}} = k dt\] integrate both sides \[-\frac{1}{x} = kt + C\] solve for x \[x = \frac{1}{C-kt}\] find C using initial value...x(0) = x_0 \[x_0 = \frac{1}{C} \rightarrow C = \frac{1}{x_0}\] after simplifying \[x(t) = \frac{x_0}{1-k x_0 t}\]

OpenStudy (anonymous):

yeah this is differential equation problems

OpenStudy (anonymous):

is this a final answer

OpenStudy (dumbcow):

i believe so

OpenStudy (dumbcow):

were you able to understand my work

OpenStudy (anonymous):

may be i understand but i confuse if it is the real answer to the question , determine the x,for all t (greater than or equal to 0)

OpenStudy (dumbcow):

x is a function of t.... so to determine x for all positive t values, we have to define the function x(t)

OpenStudy (anonymous):

ok tnx fo your help,,

OpenStudy (dumbcow):

actually just noticed something...right now the function has asymptote at t = 1/kx_0 which means x is not defined for all positive "t" values so just switch the asymptote to negative side i should have assumed the constant was negative \[C = -\frac{1}{x_0}\] this changes the function to: \[x(t) = \frac{-x_0}{1+k x_0 t}\]